$SPAD/src/input UP.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

-- Input generated from UnivariatePolynomialXmpPage
)clear all
(p,q) : UP(x,INT)
p := (3*x-1)**2 * (2*x + 8)
q := (1 - 6*x + 9*x**2)**2
p**2 + p*q
leadingCoefficient p
degree p
reductum p
gcd(p,q)
lcm(p,q)
resultant(p,q)
D p
p(2)
p(q)
q(p)
l := coefficients p
reduce(gcd,l)
content p
ux := (x**4+2*x+3)::UP(x,INT)
vectorise(ux,5)
squareTerms(p) ==
  reduce(+,[t**2 for t in monomials p])
p
squareTerms p
(r,s) : UP(a1,FRAC INT)
r := a1**2 - 2/3
s := a1 + 4
r quo s
r rem s
d := divide(r, s)
r - (d.quotient * s + d.remainder)
integrate r
integrate s
t : UP(a1,FRAC POLY INT)
t := a1**2 - a1/b2 + (b1**2-b1)/(b2+3)
u : FRAC POLY INT := t
u :: UP(b1,?)

———-

References^{[1] nothing}

4